OpenStudy (anonymous):

you have an investment of 145000.you want to invest it for 199days at a rate of 6.2%. the investment house gives you two options. they inform you that you should choose between simple interest and compound interest, calculate the difference between the 2 options.

OpenStudy (anonymous):

Are u supposed to use calculator for these or what?

OpenStudy (anonymous):

nilankshi is bored:-)

OpenStudy (nilankshi):

how u know

OpenStudy (anonymous):

It was hard to guess, being truthful..:-)

OpenStudy (anonymous):

Finance?

OpenStudy (nilankshi):

damed bored yar

OpenStudy (anonymous):

Finance is math really but there are all sorts of weird definitions, u don't know them, gets u in trouble eg 360 or 365 day year, things like that...

OpenStudy (anonymous):

I=prt and if i remember right t is in years so you will have a side calculation there is the simple interest formula

OpenStudy (anonymous):

simple = flat?

OpenStudy (anonymous):

A=p(1+r/n)^rt is compound interest formula, which again has t=time in years. Yes simple whould be flat rate return where as this will pay on previous payments

OpenStudy (anonymous):

Sorry the exponent should be nt

OpenStudy (anonymous):

formula is RxPxt/100 the othere one is A=P(1+i)N or FV= PV(1+i) powern

OpenStudy (anonymous):

Yes, basically the same... But don't u just use calculator?

OpenStudy (anonymous):

I have PV, FV, n etc on my calculator.

OpenStudy (anonymous):

Go with whatever your teacher is teaching you but those are the ones i would use